Details
-
Type:
Story
-
Status: Done
-
Resolution: Done
-
Fix Version/s: None
-
Component/s: None
-
Labels:None
-
Story Points:5
-
Sprint:TSSW Sprint - Oct 10 - Oct 24
-
Team:Telescope and Site
-
Urgent?:No
Description
Today they managed to get the CCW operating again. Also Julen tells me the Ethercat “looks promising” which is good as that was also on my SUMMIT tickets for today. I still have to slack with Jacques about the set of SUMMIT tickets I have generated.
I also downloaded the entire set of gitlab repos to make sure I have the latest. As a side effect of starting to look through them I saw they used a tool from Bosch (IndraWorks) so I am downloading that to take look. (edited)
Today there was much discussion about the OSS glycol issue, multiple paths forward are being considered.
Shawn and I participated in the pre-start inspection procedure to get some understanding of what is involved
(currently in Spanish but an english translation will be added of course). Alberto sent me the procedure he used for low-level checking of the Ethercat connections (this is another Windows based tool (from Beckhoff)). I asked Shawn to make sure that we get to keep the PC they are currently running this and other debug tools on, also downloaded a copy of it myself. We learnt that “it’ll be fine, we can use the HMI manually for now” for the Top End Chiller is no longer a viable solution as recent cable installation / cabinet reorg has made it impossible to access the HMI. I suggested we remove it from the cabinet and put it somewhere more useful (may involve ethernet through the cable wrap). Julen updated all the PXI’s but then discovered an issue somewhere in the LabVIEW code. If he gets that fixed in the next hour or so then we may try balancing the telescope….. Most folks left the mountain early to avoid potential road closures due to “anniversary of the riots”. Jacques and I did a tag up on the SUMMIT tickets but he was reluctant to consider anything beyond this week in detail. (edite
OSS discussions continued and the system was set to operate at a lower pressure as an experiment.
Other hardware issues included the capacitor bank system blowing a fuse again, Phase is doing some remote diagnosis to see if they can find out why. On the software side Julen is debugging an issue with the Encoder data not reaching the MCS. We still need to balance the telescope before it can be moved under MCS control (hopefully tomorrow). We had a 2 hour pointing-prep meeting and there will likely be another on Friday. I noticed that we do not have a lot of the Telemetry that was present during FAT (incl Oil system which I wanted to monitor during the test). I retrieved an old version of the Telemetry XML and passed it on to Russell. Tomorrow I will check with Tekniker that it is all still available from the MCS in principle.
The OSS did not fault so that is good. Also the system was started up after fuses in the cap bank were replaced, and it also stayed up. Phase investigated remotely but could find nothing wrong with the motors interaction with the cap bank. I got the MTMount CSC up using LOVE after Russell updated it. Julen has sent Telemetry doco to Russell.
The Encoder issue was also fixed by Julen (configuration problem). The rest of the Telemetry will still be available from the MCS onboard store if we need it in the meantime. The inclinometer was installed today and should be hooked up tomorrow. Balancing has been punted until tomorrow. Julen is working on the Rotator track issue that was seen in todays test (fail to track), I got the impression he knows what it is. The star-tracker camera arrived pre-mounted to its base. I think it may also get onto the telescope tomorrow.
Dave's on-going work with Tekniker, keeping track of progress and continue to push them in the right direction.